How many grams of carbon would there be in a sample that contained 8.56 x 10^49atoms?

1
Expert's answer
2021-01-28T05:13:24-0500

moles = "\frac{8.56×10⁴⁹}{6.023×10²³}" = 1.421 × 10²⁶ moles

Mass of carbon atoms = 12 × 1.421 ×10²⁶

= 17.052 × 10²⁶ g